OpenStudy (anonymous):

10. You have six $1 bills, eight $5 bills, and four $20 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill. What is P($1, then $10)? A. 77/190 B. 3/100 C. 3/95 D. 2/5

OpenStudy (anonymous):

just a question what grade r u in

OpenStudy (anonymous):

I'm in 8th

OpenStudy (anonymous):

no wonder i dont understand

OpenStudy (anonymous):

CX

OpenStudy (anonymous):

wow im usualy great at this kind of stuff

OpenStudy (anonymous):

I don't understand this either. C:

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

I don't see any $10 bills in the word problem.

OpenStudy (anonymous):

thats what got me too

OpenStudy (anonymous):

OH oops. I missed that. You have six $1 bills, eight $5 bills, two $10 bills, and four $20 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill. What is P($1, then $10)?

OpenStudy (anonymous):

Thanks for noticing

OpenStudy (anonymous):

i still dont get it and i have a math book right beside me

OpenStudy (anonymous):

Probability from one dollar bills: 3/10 Probability of 10 dollar bills: 2/19\[\frac{ 3 }{ 10 }\times \frac{ 2 }{ 19 }\]

OpenStudy (anonymous):

So, it's C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Your welcome. :)

OpenStudy (anonymous):

your welcome ^-^

OpenStudy (anonymous):

X)

OpenStudy (anonymous):

(^o-o^) ima bat

OpenStudy (anonymous):

█ ▄ ▄ o(≧o≦)o ▄ ▄ █

OpenStudy (anonymous):

I'm a DJ :)